# High-pressure floating zone crystal growth of Sr2IrO4

### v1

@@ I Introduction
For over a decade, the Mott insulating state realized in iridium oxides has garnered significant interest owing to its relevance to high-temperature superconductivity, topology, quantum computation, and other applications [1, 2, 3, 4, 5, 6, 7]. Within this family of materials, the quasi-2D layered perovskite Sr2IrO4 stands out as a close analogue in crystalline, electronic, and magnetic structure to the Mott insulator La2CuO4 [8, 9, 10, 2, 11, 7], which is a parent compound to high-temperature superconductivity upon hole doping [12, 13, 14]. While theoretical work has indeed predicted superconducting states in Sr2IrO4 at both hole-doped [15] and electron-doped [16, 17, 18, 19] compositions, definitive signatures of the superconducting state (i.e., zero resistance and diamagnetism) have thus far remained elusive. Despite this, the observation of the related cuprate phenomenology such as Fermi arcs and a pseudogap phase [2, 20, 21, 22], as well as the recent observation of a spin density wave [23] in lightly electron-doped (Sr1-xLax)2IrO4 places this system tantalizingly close to a superconducting state, demonstrating that the analogy to La2CuO4 is robust.
@@
Given these issues, there is a clear need to develop new synthesis routes which offer active control over disorder in order to understand its relationship with the observed physical properties. One possible solution is to employ the floating zone crystal growth method, which allows for well-defined control over processing parameters such as growth rate, mixing rate, heating profile, and chemical potential. The floating zone technique has been employed widely for a range of transition metal oxides, including 4d systems such as such as Sr2RuO4 [32] and Sr2RhO4 [33], but to date has not been successfully applied to any iridium oxides. This is because the growth of 4d and 5d transition metals such as Ru, Rh, and Ir from a molten phase is dramatically complicated by the tendency of the binary transition metal oxides to volatilize as a vapor phase during the growth [34, 32, 33]. For Ru and Ir in particular, this occurs via a reaction of (Ru/Ir)O2 in the presence of oxygen to form the vapor phase molecule (Ru/Ir)O3. On the other hand, the congruent melting of iridium oxides demands a high oxygen pressure because of the tendency of IrO2 to decompose into Ir metal and O2 at elevated temperatures. Further, the higher melting temperature of the iridates precludes the use of conventional halogen bulbs as heating elements, and exacerbates both of these chemical instabilities.
In the present work, we mitigate these issues using a custom-built high-pressure laser floating zone furnace [35, 36]. Relative to traditional floating zone furnaces based on halogen lamps paired with elliptical focusing mirrors, this laser-based furnace allows for finer control over the heating profile in the molten zone and makes accessible a larger range of temperatures. The furnace is capable of applying total gas pressures up to 15,000 psi to suppress material volatility, including partial pressures of oxygen of up to 3,000 psi which can be used to tune chemical potential. We demonstrate that this method allows for the first successful growth of cm3-scale single crystals of Sr2IrO4 and (Sr1-xLax)2IrO4 stabilized by moderately high oxygen partial pressures combined with relatively fast growth rates and mixing rates. In addition, we show that different processing regimes result in structural changes at the nano- and micro-scale and allow for tuning of the electronic and magnetic responses.
@@ II.1 Polycrystalline synthesis
Polycrystalline feed and seed rods were prepared by conventional solid state reaction using SrCO3 (99.99% excluding Ca or Ba, Thermo Fisher Scientific), La2O3 (99.99% metals basis, VWR International), and IrO2 (99.99% metals basis, Thermo Fisher Scientific) precursors. In order to compensate for the loss of Ir as IrO3(g) during the floating zone growth, the feed rods were prepared with an excess of Ir in the stoichiometry (Sr1-xLax)2Ir1+γO4 over the range . These reagents were ground and pressed into a pellet using a cold isostatic press under a hydrostatic pressure of psi. Samples were then reacted in air at 1100 ∘C for 18 hours. Powder X-ray diffraction collected on a ground piece of the reacted pellet reveals the monoclinic SrIrO3 phase [37] as the only secondary phase, consistent with the excess Ir stoichiometry. Pellets were then reground and pressed into a rod geometry with diameter ranging from 4 mm to 6 mm, and sintered at 1380 ∘C for 12 hours. Powder X-ray diffraction on a ground piece of the sintered rod revealed the formation of a small amount of IrO2 ( wt%).
@@ II.2 High-pressure floating zone crystal growth
Floating zone growth was carried out in a custom laser-based high-pressure furnace [35, 36]. Oxygen partial pressures ranged from psi and total pressures including an Ar overpressure ranged from psi. The volatilization of some Ir as IrO3(g) during growth was unavoidable, and resulted in the continuous deposition of solid IrO2 on the interior of the growth chamber over time as shown in Figure 1(a,b). Fortunately, the incident laser beam mitigated deposition on small areas of the quartz window surrounding the molten zone, allowing for continuous viewing of the zone even after several hours of growth.
@@ III.1 Microstructural dependence on oxygen pressure
The structural phases observed in the floating zone crystals depended heavily on the applied oxygen pressure as shown in Figure 1(f-h). At psi, insufficient Ir oxidation resulted in decomposition of the melt into Sr2IrO4, Sr9Ir3O17, and Ir metal lamellae. At the slightly higher psi, much more of the Ir metal was oxidized, thus decreasing the fraction of the Ir-poor Sr9Ir3O17 phase and increasing the fraction of the Sr2IrO4 phase. A pressure of psi was found to be sufficient to fully oxidize the Ir and avoid metal inclusions. Any local enrichment of the Ir stoichiometry beyond the Sr2IrO4 composition formed inclusions of Sr3Ir2O7 at these pressures. At a pressure of psi, evaporative losses of Ir were exacerbated by increased formation of IrO3(g), and the phase fraction of the Sr9Ir3O17 phase in the bulk increased dramatically at the expense of the Sr2IrO4 phase. These results are consistent with the phase diagram presented in Figure 2, and demonstrate that the applied oxygen pressure strongly controls the resulting phase equilibria.
Powder X-ray diffraction measurements on a ground piece of the as-grown boules fit well to a model of Sr2IrO4, with varying amounts of Sr3Ir2O7 inclusions in the bulk and Sr9Ir3O17 confined to the edges of the boule [Fig. 3(a)]. Post-growth annealing at 1400 ∘C for several hours causes the metastable Sr9Ir3O17 phase at the edges of the boule to decompose into Sr4IrO6 and Sr2IrO4, without significantly enhancing the phase fraction of Sr2IrO4 [Fig. 3(b)].
@@ III.2 Mixing-induced changes in bulk physical properties
With the ability to control the homogeneity of the melt via counter-rotation of the feed and seed rods, we demonstrate that samples grown under different mixing conditions exhibit dramatically different magnetic properties as summarized in Table 1. First we describe the behavior of the samples with maximal homogeneity in the melt, here labeled as “fast mixing”, where the counter-rotation of each rod at 50 rpm generates a total mixing rate of 100 rpm. Electrical resistivity measurements reveal insulating behavior comparable to that observed in flux-grown crystals [Fig. 4(a)] [28]. In temperature-dependent magnetization measurements, we observe the expected antiferromagnetic ordering transition near K [Fig. 4(b)]. The transition is qualitatively sharper than previously reported for flux-grown samples measured at an equivalent magnetic field [28].
In contrast, under “slow mixing” conditions, where net rotation rate was closer to 50 rpm, we observe a clear shift towards metallic behavior [Fig. 4(a)], along with a dramatic suppression of the ordering transition down to approximately K [Fig. 4(b)]. We also find that the low-temperature saturated magnetization is suppressed to about one-third of the expected value, indicating a weaker net moment in each canted AF layer [Fig. 4(c)]. Surprisingly, the measured temperature-dependent magnetization in the intermediate range of appears to reach the full saturated magnetization (), despite the relatively weak applied field of . This is potentially because the planar moments are pinned by disorder to a more polarized state during the (Z)FCW measurements. Further, a second temperature scale appears at whereupon the magnetization decreases sharply away from the fully saturated value and converges towards a magnetization comparable to that of the sample grown under fast mixing conditions – suggesting a reentrance of the staggered “” phase at low temperature.
@@
The longitudinal resistivity of the “fast mixing” samples shows slightly more metallic behavior than flux-grown crystals, likely due to percolation through inclusions of secondary phases embedded in the parent Sr2IrO4 matrix [Fig. 4(a)]. Consistent with the observed magnetization data, samples grown with slow mixing are significantly more metallic than samples grown with fast mixing, with a resistivity comparable to that of electron-doped samples of (Sr1-xLax)2IrO4 with [28].
@@
In order to further verify the above interpretation, we introduced a small fraction of La dopants in the amount of during a slow mixing growth, which, in the case of electron (hole) doping should cause further suppression (recovery) of the electrical resistivity and magnetic ordering. We find that this La-doped sample exhibits a stark recovery of the strongly insulating behavior observed for fast mixed samples [Fig. 4(a)], strongly supporting the notion of compensation for defect-induced, hole-like carriers. We also find a remarkable recovery of the magnetic order, with a sharp ordering transition observed in the temperature-dependent magnetization and a significant increase in the saturated magnetization [Fig. 4(b,c)]. A Curie-Weiss fit to the inverse magnetic susceptibility for this sample yields a Curie-Weiss temperature of and an effective paramagnetic moment of . Additionally, it is notable that the irreversibility observed in this lightly-doped sample is smaller than both the fast-mixed and the slow-mixed Sr2IrO4, which may indicate light hole doping in the fast-mixed samples as well. Taken together, these findings strongly suggest that slow mixing enables the synthesis of vacancy-doped Sr2Ir1-yO4, and that substitution with La3+ compensates the holes generated by these vacancies with electrons to induce a recovery of the insulating state and magnetic ordering. If true, then a future growth trial at the same composition of under fast mixing conditions should exhibit enhanced suppression of the magnetic ordering relative to its slow-mixed counterpart.
@@ IV Conclusions
We have successfully demonstrated the first floating zone crystal growth of an iridate compound via the growth of Sr2IrO4, enabled via high pressure. To our knowledge, these samples are the largest (cm3-scale) crystals currently available. A moderate O2 pressure stabilizes the correct oxidation state, and an inert overpressure of Ar helps mitigate IrO3 volatility during growth. We demonstrate that active control over mixing during the growth allows for tuning of the electronic and magnetic response of the material across a dramatic range of properties, likely due to the introduction of Ir vacancies which introduce holes into the system. The successful realization of these large-scale crystals may be expected to allow for deeper investigations of the strongly spin-orbit coupled physics in Sr2IrO4 by weakly interacting probes such as neutron scattering. Further, this new regime of growth conditions may be amenable to controlling both disorder and solubility of dopants which may be an important tuning parameter in the pursuit of unconventional superconductivity in this member of the iridate family of compounds. Lastly, we stress that high-pressure oxygen appears to be an important tuning parameter here, and hope that these breakthroughs translate to other iridium oxide-based compounds of interest.

### v2

@@ I Introduction
For over a decade, the Mott insulating state realized in iridium oxides has garnered significant interest owing to its relevance to high-temperature superconductivity, topology, quantum computation, and other applications [1, 2, 3, 4, 5, 6, 7]. Within this family of materials, the quasi-2D layered perovskite Sr2IrO4 stands out as a close analogue in crystalline, electronic, and magnetic structure to the Mott insulator La2CuO4 [8, 9, 10, 2, 11, 7], which is a parent compound to high-temperature superconductivity upon hole doping [12, 13, 14]. While theoretical work has indeed predicted superconducting states in Sr2IrO4 at both hole-doped [15] and electron-doped [16, 17, 18, 19] compositions, definitive signatures of the superconducting state (i.e., zero resistance and diamagnetism) have thus far remained elusive. Despite this, the observation of the related cuprate phenomenology such as Fermi arcs and a pseudogap phase [2, 20, 21, 22], as well as the recent observation of a spin density wave [23] in lightly electron-doped (Sr1-xLax)2IrO4, places this system tantalizingly close to a superconducting state, demonstrating that the analogy to La2CuO4 is robust.
@@
Given these issues, there is a clear need to develop new synthesis routes which offer active control over disorder in order to understand its relationship with the observed physical properties. One possible solution is to employ the floating zone crystal growth method, which allows for well-defined control over processing parameters such as growth rate, mixing rate, heating profile, and chemical potential. The floating zone technique has been employed widely for a range of transition metal oxides, including 4d systems such as such as Sr2RuO4 [32] and Sr2RhO4 [33], but to date has not been successfully applied to any iridium oxides. This is because the growth of 4d and 5d transition metals such as Ru, Rh, and Ir from a molten phase is dramatically complicated by the tendency of the binary transition metal oxides to volatilize as a vapor phase during the growth [34, 32, 33]. For Ru and Ir in particular, this occurs via a reaction of (Ru/Ir)O2 in the presence of oxygen to form the vapor phase molecule (Ru/Ir)O3. On the other hand, the congruent melting of iridium oxides demands a high oxygen pressure because of the tendency of IrO2 to decompose into Ir metal and O2 at elevated temperatures. Further, the higher melting temperature of the iridates precludes the use of conventional halogen bulbs as heating elements and exacerbates both of these chemical instabilities.
In the present work, we mitigate these issues using a custom-built high-pressure laser floating zone furnace [35, 36]. Relative to traditional floating zone furnaces based on halogen lamps paired with elliptical focusing mirrors, this laser-based furnace allows for finer control over the heating profile in the molten zone and makes accessible a larger range of temperatures. The furnace is capable of applying total gas pressures up to 15,000 psi (1,000 bar) to suppress material volatility, including partial pressures of oxygen of up to 3,000 psi (200 bar) which can be used to tune chemical potential. We demonstrate that this method allows for the first successful growth of cm3-scale single crystals of Sr2IrO4 and (Sr1-xLax)2IrO4 stabilized by moderately high oxygen partial pressures combined with relatively fast growth rates and mixing rates. In addition, we show that different processing regimes result in structural changes at the nano- and micro-scale and allow for tuning of the electronic and magnetic responses.
@@ II.1 Polycrystalline synthesis
Polycrystalline feed and seed rods were prepared by conventional solid state reaction using SrCO3 (99.99% excluding Ca or Ba, Thermo Fisher Scientific), La2O3 (99.99% metals basis, VWR International), and IrO2 (99.99% metals basis, Thermo Fisher Scientific) precursors. In order to compensate for the loss of Ir as IrO3(g) during the floating zone growth, the feed rods were prepared with an excess of Ir in the stoichiometry (Sr1-xLax)2Ir1+γO4 over the range . These reagents were ground and pressed into a pellet using a cold isostatic press under a hydrostatic pressure of psi (3275 bar). Samples were then reacted in air at 1100 ∘C for 18 hours. Powder X-ray diffraction collected on a ground piece of the reacted pellet reveals the monoclinic SrIrO3 phase [37] as the only secondary phase, consistent with the excess Ir stoichiometry. Pellets were then reground and pressed into a rod geometry with diameter ranging from 4 mm to 6 mm, and sintered at 1380 ∘C for 12 hours. Powder X-ray diffraction on a ground piece of the sintered rod revealed the formation of a small amount of IrO2 ( wt%).
@@ II.2 High-pressure floating zone crystal growth
Floating zone growth was carried out in a custom laser-based high-pressure furnace [35, 36]. Oxygen partial pressures ranged from psi (140 bar) and total pressures including an Ar overpressure ranged from psi ( bar). The volatilization of some Ir as IrO3(g) during growth was unavoidable, and resulted in the continuous deposition of solid IrO2 on the interior of the growth chamber over time as shown in Figure 1(a,b). Fortunately, the incident laser beam mitigated deposition on small areas of the quartz window surrounding the molten zone, allowing for continuous viewing of the zone even after several hours of growth.
@@ III.1 Microstructural dependence on oxygen pressure
The structural phases observed in the floating zone crystals depended heavily on the applied oxygen pressure as shown in Figure 1(f-h). At psi (14 bar), insufficient Ir oxidation resulted in decomposition of the melt into Sr2IrO4, Sr9Ir3O17, and Ir metal lamellae. At the slightly higher psi, much more of the Ir metal was oxidized, thus decreasing the fraction of the Ir-poor Sr9Ir3O17 phase and increasing the fraction of the Sr2IrO4 phase. A pressure of psi (70 bar) was found to be sufficient to fully oxidize the Ir and avoid metal inclusions. Any local enrichment of the Ir stoichiometry beyond the Sr2IrO4 composition formed inclusions of Sr3Ir2O7 at these pressures. At a pressure of psi (140 bar), evaporative losses of Ir were exacerbated by increased formation of IrO3(g), and the phase fraction of the Sr9Ir3O17 phase in the bulk increased dramatically at the expense of the Sr2IrO4 phase. These results are consistent with the phase diagram presented in Figure 2, and demonstrate that the applied oxygen pressure strongly controls the resulting phase equilibria.
For the remainder of the present analysis, we focus on samples which were grown under the optimized 1,000 psi (70 bar) oxygen partial pressure. Powder X-ray diffraction measurements on a ground piece of the as-grown boules fit well to a model of Sr2IrO4, with varying amounts of Sr3Ir2O7 inclusions in the bulk and Sr9Ir3O17 confined to the edges of the boule [Fig. 3(a)]. We note that we are not able to distinguish between the lower-symmetry model reported for crystals grown in Pt crucibles [40, 41, 42, 43] and the model recently reported for crystals grown in Ir crucibles [26]. Post-growth annealing at 1400 ∘C for several hours causes the metastable Sr9Ir3O17 phase at the edges of the boule to decompose into Sr4IrO6 and Sr2IrO4, without significantly enhancing the phase fraction of Sr2IrO4 [Fig. 3(b)].
@@ III.2 Mixing-induced changes in bulk physical properties
With the ability to control the homogeneity of the melt via counter-rotation of the feed and seed rods, we demonstrate that samples grown under different mixing conditions exhibit dramatically different magnetic properties as summarized in Table 1. First we describe the behavior of the samples with maximal homogeneity in the melt, here labeled as “fast mixing”, where the counter-rotation of each rod at 50 rpm generates a total mixing rate of 100 rpm. Electrical resistivity measurements reveal insulating behavior comparable to that observed in flux-grown crystals [Fig. 4(a)] [28]. In temperature-dependent magnetization measurements, we observe the expected antiferromagnetic ordering transition near K [Fig. 4(b)]. The slightly lower resistivity and in the floating zone crystals may arise from a small fraction of dopants (i.e., a slight off-stoichiometry below the level detectable via energy-dispersive spectroscopy or XRD measurements), though these properties are also expected to be sensitive to other forms of disorder such as isoelectronic substitution [44]. Despite this, the magnetic ordering transition is sharper that previously reported for flux-grown samples measured at an equivalent magnetic field [28], suggesting a globally cleaner magnetic state even in the presence of a weak doping effect. The likely cause here is a subtle Ir off-stoichiometry due to the evaporative loss of Ir that mimics the Pt impurities discussed by Kim et al. [26]. Further optimization of the floating zone growth can likely allow this small off-stoichiometry to be mitigated or it may be compensated via La doping.
In contrast, under “slow mixing” conditions, where the net rotation rate was closer to 50 rpm, we observe a clear shift towards less insulating behavior [Fig. 4(a)], along with a dramatic suppression of the ordering transition down to approximately K [Fig. 4(b)]. We also find that the low-temperature saturated magnetization is suppressed to about one-third of the expected value, indicating a weaker net moment in each canted AF layer [Fig. 4(c)]. Surprisingly, the measured temperature-dependent magnetization in the intermediate range of appears to reach the full saturated magnetization (), despite the relatively weak applied field of . This is potentially because the planar moments are pinned by disorder to a more polarized state during the (Z)FCW measurements. Further, a second temperature scale appears at whereupon the magnetization decreases sharply away from the fully saturated value and converges towards a magnetization comparable to that of the sample grown under fast mixing conditions – suggesting a reentrance of the staggered “” phase at low temperature.
@@
The longitudinal resistivity of the “fast mixing” samples shows slightly less insulating behavior than flux-grown crystals likely due to percolation through inclusions of secondary phases embedded in the parent Sr2IrO4 matrix [Fig. 4(a)]. Consistent with the observed magnetization data, samples grown with slow mixing are significantly less insulating than samples grown with fast mixing, with a resistivity comparable to that of electron-doped samples of (Sr1-xLax)2IrO4 with [28]. Fitting the high-temperature resistivity of the fast mixing sample to an activation law yields an energy gap , which is comparable to the bulk gap measured for flux-grown crystals (Fig. S2). We emphasize here that the gap extracted from these fits is not representative of the true Mott gap, but is instead sensitive to in-gap states which may originate from the presence of dopants or defects in the bulk Sr2IrO4 sample used for transport measurements.
@@
In order to further verify the above interpretation, we introduced a small fraction of La dopants in the amount of (determined via energy-dispersive spectroscopy measurements) during a slow mixing growth, which, in the case of electron (hole) doping should cause further suppression (recovery) of the electrical resistivity and magnetic ordering. We find that this La-doped sample exhibits a stark recovery of the strongly insulating behavior observed for fast mixed samples [Fig. 4(a)], strongly supporting the notion of compensation for defect-induced, hole-like carriers. We also find a remarkable recovery of the magnetic order, with a sharp ordering transition observed in the temperature-dependent magnetization and a significant increase in the saturated magnetization [Fig. 4(b,c)]. A Curie-Weiss fit to the inverse magnetic susceptibility for this sample yields a Curie-Weiss temperature of and an effective paramagnetic moment of . Additionally, it is notable that the irreversibility observed in this lightly-doped sample is smaller than both the fast-mixed and the slow-mixed Sr2IrO4, which may indicate light hole doping in the fast-mixed samples as well. Taken together, these findings strongly suggest that slow mixing enables the synthesis of vacancy-doped Sr2Ir1-yO4, and that substitution with La3+ compensates the holes generated by these vacancies with electrons to induce a recovery of the insulating state and magnetic ordering. If true, then a future growth trial at the same composition of under fast mixing conditions should exhibit enhanced suppression of the magnetic ordering relative to its slow-mixed counterpart.
@@ IV Conclusions
We have successfully demonstrated the first floating zone crystal growth of an iridate compound via the growth of Sr2IrO4, enabled via high pressure. To our knowledge, these samples are the largest (cm3-scale) crystals currently available. A moderate O2 pressure stabilizes the correct oxidation state, and an inert overpressure of Ar helps mitigate IrO3 volatility during growth. We demonstrate that active control over mixing during the growth allows for tuning of the electronic and magnetic response of the material across a dramatic range of properties, likely due to the formation of Ir vacancies which introduce holes into the system. The successful realization of these large-scale crystals may be expected to allow for deeper investigations of the strongly spin-orbit coupled physics in Sr2IrO4 by weakly interacting probes such as neutron scattering. Further, this new regime of growth conditions may be amenable to controlling both disorder and solubility of dopants which may be an important tuning parameter in the pursuit of unconventional superconductivity in this member of the iridate family of compounds. Lastly, we stress that high-pressure oxygen appears to be an important tuning parameter here, and hope that these breakthroughs translate to other iridium oxide-based compounds of interest.
@@ VII Data availability
Supplemental Information
| Spectrum | Element | Line Type | Weight % | Atomic % | Sr:Ir (m/m) |
| (a) | Sr | L series | 16.89 | 2.64 | |
| Ir | L series | 6.40 | |||
| (b) | Sr | L series | 16.15 | 2.07 | |
| Ir | L series | 7.82 | |||
| (c) | Sr | L series | 17.55 | 2.00 | |
| Ir | L series | 8.79 | |||
| (d) | Sr | L series | 12.60 | 2.48 | |
| Ir | L series | 5.07 | |||
| (e) | Sr | L series | 12.51 | 1.98 | |
| Ir | L series | 6.32 | |||
| (f) | Sr | L series | 10.58 | 1.79 | |
| Ir | L series | 5.91 |